OpenStudy (anonymous):

what is the sq. root of 192?

OpenStudy (saifoo.khan):

\[8\sqrt{3}\]

OpenStudy (anonymous):

i just have a question... isnt a sq. root a number squared? if so, how is it \[8\sqrt{3}\]

OpenStudy (anonymous):

192 = 64*3 ===> sqrt(192) = 8*sqrt(3)

OpenStudy (saifoo.khan):